# citations.bib

### v1

@@ III-D RoBERTa
Similarly to the LSTM models built using the BERT embeddings, the LSTM model with the best performance was achieved using the Adam optimizer with an initial learning rate of 0.0001. Table  IX highlights that neural networks created using RoBERTa embeddings consistently achieved better performance than those trained using BERT embeddings, with the highest performing model achieving an accuracy of 87%.
TABLE IX: RoBERTa Metrics
Hyper-
parameters
@@
Epoch Loss Accuracy Precision Recall
F1-
Score
Adam
-LR: 0.0001
46 0.3152 87.0% 78.8% 89.4% 83.8%
SGD
-LR: 0.0001
-Mom: 0.01
45 0.3385 84.2% 73.9% 89.4% 80.9%
SGD
-LR: 0.0001
-Mom: 0.001
32 0.3737 83.3% 73.1% 87.6% 79.7%
SGD
-LR: 0.01
-Mom: 0.01
29 0.3489 83.0% 73.1% 86.6% 79.3%
SGD
-LR: 0.0001
-Mom: 0.0001
39 0.3721 82.9% 72.8% 86.8% 79.2%
@@ IV Discussion
Neural networks created with the word2vec: CBOW embeddings achieved the best performance with a peak accuracy of 92.3% when the Adam optimizer was used with a learning rate of 0.0001. Similarly, models utilizing word2vec: Skip-Gram embeddings also demonstrated strong performance, achieving an accuracy of 90.6% with the same optimizer and learning rate. However, neither word2vec: CBOW and Skip-Gram performed well when the SGD optimizer was used with a lower learning rate (0.0001). Nevertheless, the behavior displayed by the word2vec CBOW accuracy and loss graphs indicate that additional epochs may have resulted in better performance.
The higher performance of word2vec CBOW over Skip-Gram is somewhat interesting. Once preprocessing was completed, the number of unique tokens within the LLVM functions was somewhat limited. This was due to the way the preprocessing limited unnecessary uniqueness within tokenized LLVM functions. The CBOW model’s technique tends to produce more accurate embeddings for representing tokens with higher frequency than the Skip-Gram model [mikolov2013efficient]. However, Skip-Gram models typically produce more accurate embeddings than CBOW models when trained with smaller datasets [mikolov2013efficient]. Thus, in this research, the accuracy gains achieved through token frequency resulted in better CBOW performance despite the relatively small training dataset.
The top accuracy of neural networks trained using BERT and RoBERTa embeddings were not as high as those trained using word2vec embeddings. This implies that, for smaller datasets, the simpler techniques used by word2vec outperform the more complex context-aware techniques used by BERT and RoBERTa. This result is not overly surprising considering the BERT model was pretrained using the BooksCorpus (800M words) and English Wikipedia (2,500M words) [devlin2019bert] and RoBERTa was trained on ten times that amount of data [liu2019roberta]. Thus, it is suspected that neural networks built using embeddings from BERT and RoBERTa models trained on larger datasets would likely achieve higher performance. However, despite the smaller training dataset used to train the BERT and RoBERTa models, it is interesting to note that neural networks utilizing these embeddings, optimized with SGD and a learning rate of 0.0001, outperformed the word2vec models trained under the same conditions. Furthermore, these models achieved a relatively strong top performance, achieving high accuracy rates of 87.0% with RoBERTa and 84.1% with BERT when the Adam optimizer was used with a learning rate of 0.0001. Thus, even though these models are meant to be trained with large amounts of data, they still manage to perform moderately well with smaller datasets.
@@ V Related Work
Few papers used a common NLP embedding model other than word2vec in their research. [RN67] used RoBERTa within their process to learn semantic relationships within assembly code and passed these embeddings to a Message Passing Neural Network (MPNN) for vulnerability detection. [gallagherllvm] custom-built a RoBERTa model with LLVM code to learn the code-level semantics to detect vulnerable code. [han2022binary] trained a BERT model with pcode and utilized the generated embeddings to train machine learning models to differentiate between vulnerable and non-vulnerable code. Finally, [koo2021semantic] employed BERT and [10.1145/3460120.3484587] used a GPT model to create embeddings for code similarity detection.
Recurrent neural networks commonly used in natural language processing are regularly used to detect vulnerabilities in compiled binaries. For example, [RN36, RN40, RN10, RN9, RN57, RN47, RN51, RN58, schaad2022deeplearningbased] used an LSTM neural network and [RN9, RN42, RN57, schaad2022deeplearningbased] applied a GRU neural network for vulnerability detection. Furthermore, [RN7, RN10, RN9, RN57, RN45, RN51, RN55, schaad2022deeplearningbased] used a BLSTM, and [RN10, RN9, RN57, RN45, RN46, RN51, RN55, schaad2022deeplearningbased] used a BGRU neural network as part of the research they performed.
@@ VI Future Research
One of the most significant limitations encountered during the current study was the low number of training samples. Once duplicate functions were removed, the number of functions available to train embedding models and LSTM neural networks decreased drastically. This could indicate that the preprocessing phase may have resulted in many samples with slight code nuances being flattened to a single form. However, this could also suggest that the number of functions that did nothing other than call secondary functions comprised a much larger portion of the dataset than originally expected. Future research will re-evaluate the preprocessing performed on each function to determine how much preprocessing resulted in the sample decrease.
It is postulated that the BERT and RoBERTa embeddings may perform better with more data, so significant emphasis in future research will be on identifying ways to increase the number of quality samples used to train the embedding models. Furthermore, it is possible that additional vulnerable and non-vulnerable code samples could also affect LSTM neural network performance. Thus, a second emphasis for future research would be to identify ways to increase the number of code samples that could be used to train these models.
@@
Finally, the study [arp2022and] states that data snooping is one of the biggest pitfalls in current research on the application of machine learning to cyber security. Data snooping occurs when training and validation data are not properly separated. One may argue that data snooping is present within the current paper because all code samples, including the samples used to train the neural networks, were used to train the embedding models. However, it should be noted that strict data segmentation was used when training each neural network. However, the effects of using these data samples when training the embedding models are unknown and will be explored further in later research.
@@ VII Conclusion
The most significant contribution of the current research is that it is the first study to explore the use of multiple bidirectional transformer (i.e. BERT & RoBERTa) embeddings built using LLVM code to train neural networks to detect vulnerabilities in the lifted code. In addition, it explored the effects of Adam and SGD optimizers, and compared the performance of models using varied values for momentum. This study also identified vital limitations that more complex NLP embeddings introduce to the process of using embeddings for vulnerability discovery in compiled binaries, in contrast to simpler word2vec embeddings. For example, increased embedding max_lengths results in significant memory consumption that limits the maximum number of tokens that can be processed in any code block.
At the end of the project, it was determined that, when using embeddings generated from NLP models trained using a limited number of samples (i.e., 118K), the LSTM built using word2vec: CBOW embeddings with the Adam optimizer and an initial learning rate of 0.0001 achieved the highest accuracy (92.3%). Furthermore, although LSTMs trained using word2vec: Skip-Gram embeddings struggled to achieve accuracy gains with an SGD optimizer and lower learning rates, it achieved a 90.6% accuracy when the Adam optimizer was used with the same initial learning rate. This accuracy was higher than the highest accuracy that was achieved by any LSTM neural network trained using BERT or RoBERTa embeddings. The results suggest that complex contextual NLP embeddings may not provide advantages over simpler word2vec models for this task when these models are trained using smaller datasets. The comparative results provide novel insights into selecting optimal embeddings for learning compiler-independent semantic code representations to advance machine learning detection of vulnerabilities in compiled code.

### v2

@@ III-C RoBERTa Metrics
The LSTM model with the best performance was achieved using the SGD optimizer with a learning rate of 0.0001 and a momentum of 0.001. Table  VI highlights that neural networks created using RoBERTa embeddings consistently achieved better performance than those trained using BERT embeddings, with the highest performing model achieving an accuracy of 88.8%.
TABLE VI: RoBERTa Metrics
@@
Hyper- parameters Epoch Loss Accuracy Precision Recall F1- Score
SGD -LR: 0.01 -Mom: 0.01 31 0.2625 86.6% 82.2% 81.0% 81.6%
SGD -LR: 0.0001 -Mom: 0.01 47 0.2454 88.7% 79.0% 94.3% 85.9%
SGD -LR: 0.0001 -Mom: 0.001 45 0.2708 88.8% 87.6% 81.0% 84.2%
SGD -LR: 0.0001 -Mom: 0.0001 49 0.2868 88.6% 84.0% 84.9% 84.5%
III-D BERT Metrics
The LSTM model that performed best with BERT embeddings was built using the SGD optimizer with a learning rate of 0.01. Table  VII highlights that the best-performing LSTM model achieved an accuracy of 88.6% with an F1-Score of 85.2%.
TABLE VII: BERT Metrics
Hyper- parameters Epoch Loss Accuracy Precision Recall F1- Score
SGD -LR: 0.0001 -Mom: 0.01 20 0.2913 88.6% 81.0% 90.0% 85.2%
@@ IV Discussion
As shown in Table VIII, neural networks built with word2vec Skip-Gram embeddings achieved the highest performance, reaching a peak accuracy of 92.0% when using the Adam optimizer with a learning rate of 0.001. Similarly, models utilizing word2vec: CBOW embeddings also demonstrated relatively strong performance, achieving a high accuracy of 87.5% with the same optimizer and learning rate. However, neither word2vec: CBOW and Skip-Gram performed well when the SGD optimizer was used with a lower learning rate (0.0001). Nevertheless, the behavior displayed by the word2vec CBOW loss graphs indicates that additional epochs may have resulted in better performance.
TABLE VIII: Overall LSTM Model Comparison
Hyper- parameters Epoch Loss Accuracy Precision Recall F1- Score
word2vec Skip-Gram Adam -LR: 0.001 43 0.1848 92.0% 85.2% 94.6% 89.6%
RoBERTa SGD -LR: 0.0001 -Mom: 0.001 45 0.2708 88.8% 87.6% 81.0% 84.2%
word2vec CBOW Adam -LR: 0.001 34 0.2528 87.5% 78.0% 91.8% 84.3%
BERT SGD -LR: 0.0001 -Mom: 0.01 20 0.2913 88.6% 81.0% 90.0% 85.2%
The higher performance of word2vec Skip-Gram over CBOW is somewhat interesting. After preprocessing, the number of unique tokens within the LLVM functions was limited, with all functions represented by only 332 unique tokens. This was due, in part, to the way that preprocessing limited unnecessary uniqueness within tokenized LLVM functions. The CBOW model’s technique tends to produce more accurate embeddings for representing tokens with higher frequency than the Skip-Gram model [mikolov2013efficient]. However, Skip-Gram models typically produce more accurate embeddings than CBOW models when trained with smaller datasets [mikolov2013efficient]. Thus, in this research, Skip-Grams’ capabilities to represent token-level relationships using datasets of limited size resulted in better performance despite CBOW’s accuracy gains achieved through token frequency.
The top accuracy of neural networks trained using BERT and RoBERTa embeddings was lower than the model trained using word2vec Skip-Gram embeddings. This implies that, for smaller datasets, the simpler techniques used by word2vec Skip-Gram outperform the more complex context-aware techniques used by BERT and RoBERTa. This result is not overly surprising considering that the original BERT model was pre-trained using BooksCorpus (800M words) and English Wikipedia (2,500M words) [devlin2019bert] and RoBERTa was trained on ten times that amount of data [liu2019roberta]. Thus, it is suspected that neural networks built using embeddings from BERT and RoBERTa models trained on larger datasets would likely achieve higher performance. However, despite the smaller training dataset used to train the BERT and RoBERTa models, it is interesting to note that neural networks utilizing these embeddings, optimized with SGD and a learning rate of 0.0001, outperformed the word2vec models trained under the same conditions. Furthermore, these models achieved relatively strong top performance, achieving high accuracy rates of 88.8% with RoBERTa and 88.6% with BERT. Thus, even though these models are meant to be trained with large amounts of data, they still manage to perform moderately well with smaller datasets.
@@ V Related Work
Few papers used a common NLP embedding model other than word2vec in their research. [RN67] used RoBERTa within their process to learn semantic relationships within assembly code and passed these embeddings to a Message Passing Neural Network (MPNN) for vulnerability detection. [gallagherllvm] custom-built a RoBERTa model with LLVM code to learn the code-level semantics to detect vulnerable code. [han2022binary] trained a BERT model with pcode and utilized the generated embeddings to train machine learning models to differentiate between vulnerable and non-vulnerable code. Finally, [koo2021semantic] used BERT, and [10.1145/3460120.3484587] used a GPT model to create embeddings for code similarity detection.
Recurrent neural networks commonly used in NLP are regularly used to detect vulnerabilities in compiled binaries. For example, [RN36, RN40, RN10, RN9, RN57, RN47, RN51, RN58, schaad2022deeplearningbased] used an LSTM neural network and [RN9, RN42, RN57, schaad2022deeplearningbased] applied a GRU neural network for vulnerability detection. Furthermore, [RN7, RN10, RN9, RN57, RN45, RN51, RN55, schaad2022deeplearningbased] used a BLSTM, and [RN10, RN9, RN57, RN45, RN46, RN51, RN55, schaad2022deeplearningbased] used a BGRU neural network as part of the research they performed.
@@ VI Future Research
It is postulated that the BERT and RoBERTa embeddings may perform better with more data, so significant emphasis in future research will be on identifying ways to increase the number of quality samples used to train the embedding models. Furthermore, it is possible that additional vulnerable and non-vulnerable code samples could also affect LSTM neural network performance. Thus, a second emphasis for future research would be identifying ways to increase the number of code samples that could be used to train these models.
@@
Finally, the study [arp2022and] states that data snooping is one of the biggest pitfalls in current research on the application of machine learning to cyber security. Data snooping occurs when training and validation data are not properly separated. The current study does not suffer from data snooping since the samples within the dataset used to train the LSTM models were never used to train any of the LLM models. However, due to the prominence of data snooping in the current literature [arp2022and], it would be worthwhile to perform additional experiments to understand just how significant the effect of data snooping would be on current results.
@@ VII Conclusion
The most significant contribution of the current research is that it is the first study to explore the use of multiple bidirectional transformers (i.e. BERT & RoBERTa) embeddings built using LLVM code to train neural networks to detect vulnerabilities in the lifted code. In addition, it explored the effects of Adam and SGD optimizers, and compared the performance of models using varied values for momentum. This study also identified vital limitations that more complex NLP embeddings introduce to the process of using embeddings for vulnerability discovery in compiled binaries, in contrast to simpler word2vec embeddings. For example, increased embedding max_lengths results in significant memory consumption that limits the maximum number of tokens that can be processed in any code block.
At the end of the project, it was determined that, when using embeddings generated from NLP models trained using a limited number of samples (i.e., 48K), the LSTM built using word2vec: Skip-Gram embeddings with the Adam optimizer and an initial learning rate of 0.001 achieved the highest accuracy (92.0%). However, this top accuracy is only slightly better than those of RoBERTa (88.8%) and BERT (88.6%). The results suggest that complex contextual NLP embeddings may not provide advantages over simpler word2vec models for this task when these models are trained using smaller datasets. The comparative results provide novel insights into selecting optimal embeddings for learning compiler-independent semantic code representations to advance machine learning detection of vulnerabilities in compiled code.